OpenStudy (anonymous):

A 3.6 kg object is accelerated from rest to a speed of 22.9 m/s in 34 s.What average force was exerted on the object during this period of acceleration? Answer in units of N

OpenStudy (anonymous):

You can use the equation Vf=Vi+at in order to find the acceleration. Vf is final velocity, or 22.9, and Vi is initial velocity, which is zero, so the equation is Vf=at. Divide by t in order to solve for acceleration. a=Vf/t You are given Vf and t, so plug them in. a=22.9/34 a=.674 m/s^2 This should be the average acceleration over the 34 seconds. Now, in order to find force, use Newton's second law, F=ma. F=ma F=3.6*.674 F=2.425 N Hope that helps!

OpenStudy (anonymous):

thank you poeticalto